Solve the following equations and check your results. 3x = 2x + 18 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3x=2x+18
⇒3x-2x=18
⇒x=18
PuttingthevalueofxinRHSandLHSweget,
3×18=(2×18)+18
⇒54=54
⇒LHS=RHS

Thanks